OpenStudy (idek-ok):

How do I cross multiply -2/4 by 2/sqrt(2)?

MARC (marc_d):

\[-\frac{ 2 }{ 4 }=\frac{ 2 }{ \sqrt{2} }\]like this?

OpenStudy (idek-ok):

-2/4 * 2/sqrt(2)

MARC (marc_d):

\[\frac{ 2 }{ 4 }\times \frac{ 2 }{ \sqrt{2} }\]

OpenStudy (idek-ok):

Yes

MARC (marc_d):

alrite

OpenStudy (idek-ok):

I'm sorry if I got the terms wrong. This isn't my strongest point in math, and I'm struggling with this concept - moreso since I'm currently working on the unit circle

OpenStudy (idek-ok):

So an explanation would really be helpful

MARC (marc_d):

it's okay.. step 1:simlifyit it until it is in simplest form\[-\frac{ 2 }{ 4 }\times \frac{ 2 }{ \sqrt{2} }\]

MARC (marc_d):

*simplify

OpenStudy (idek-ok):

How do I simplify the 2 and the square root of 2? or am i completely off .-.

MARC (marc_d):

\[-\frac{ 1\cancel{2} }{ 2\cancel{4} }\times \frac{ 2 }{ \sqrt{2} }\]

OpenStudy (idek-ok):

okay…

OpenStudy (idek-ok):

So 1/sqrt(2) ?

MARC (marc_d):

then,s\[-\frac{ 1 }{ 1\cancel{2} }\times \frac{ 1\cancel{2} }{ \sqrt{2} }\]

OpenStudy (idek-ok):

Okay, I see where that comes from! Cool!

MARC (marc_d):

it should hv the negative sign.. :)

OpenStudy (idek-ok):

*-1/sqrt(2) … sorry, figured haha

OpenStudy (idek-ok):

So the only option I see for the answer is -sqrt (2)

OpenStudy (idek-ok):

Is there another step to that?

MARC (marc_d):

yep

MARC (marc_d):

hint:\[a^-{\frac{ 1 }{ 2 }}=\frac{ 1 }{ a^{\frac{ 1 }{ 2 }} }\]

OpenStudy (idek-ok):

so would the -1 / sqrt(2) become -sqrt(2) according to that then?

OpenStudy (idek-ok):

or am i totally off

MARC (marc_d):

lt will be like this \[-\frac{ 1 }{ \sqrt{2} }\]\[=\frac{ 1 }{ 2^{\frac{ 1 }{ 2 }} }\]

MARC (marc_d):

yep,that's right!

OpenStudy (idek-ok):

WOOOOOO YEA

OpenStudy (idek-ok):

Thank you so much! :DDD I really appreciate it

MARC (marc_d):

yw! :)